8. The XY Co. has sales of $198,000, net profit of $17,500, fixed assets of $125,400, and current assets of $44,300. What is the total asset turnover rate?
Answer #1

Answer to the question:

Total Asset Turnover ratio = Total Turnover or revenue/Average total assets

= 198000 / (125400+44300)/2

= 198000 / 84850

=2.33 times

Note: It is assumed that opening assets are zero, hence average assets = (0 + closing value)/2

IF we assume that opening assets are same as that of closing then

Total Asset Turnover ratio = Total Turnover or revenue/ Total assets

= 198000 / (125400+44300)

= 1.16676 times